# EditSplat: Multi-View Fusion and Attention-Guided Optimization for View-Consistent 3D Scene Editing with 3D Gaussian Splatting

### v1

@@ Abstract
Recent advancements in 3D editing have highlighted the potential of text-driven methods in real-time, user-friendly AR/VR applications. However, current methods rely on 2D diffusion models without adequately considering multi-view information, resulting in multi-view inconsistency. While 3D Gaussian Splatting (3DGS) significantly improves rendering quality and speed, its 3D editing process encounters difficulties with inefficient optimization, as pre-trained Gaussians retain excessive source information, hindering optimization. To address these limitations, we propose EditSplat, a novel 3D editing framework that integrates Multi-view Fusion Guidance (MFG) and Attention-Guided Trimming (AGT). Our MFG ensures multi-view consistency by incorporating essential multi-view information into the diffusion process, leveraging classifier-free guidance from the text-to-image diffusion model and the geometric properties of 3DGS. Additionally, our AGT leverages the explicit representation of 3DGS to selectively prune and optimize 3D Gaussians, enhancing optimization efficiency and enabling precise, semantically rich local edits. Through extensive qualitative and quantitative evaluations, EditSplat achieves superior multi-view consistency and editing quality over existing methods, significantly enhancing overall efficiency. Project website: https://kuai-lab.github.io/editsplat2024/
@@
Figure 1: Result of EditSplat. EditSplat demonstrates its capability for flexible and high-quality 3D scene editing when applied to a pre-trained 3D Gaussian Splatting model guided by text instructions. Through its design focused on multi-view consistency, efficient optimization, and precise semantic local editing, EditSplat produces fine-grained, coherent multi-view renderings.
@@ 1 Introduction
InstructNeRF2NeRF (IN2N) [12] introduced the pipeline for text-driven 3D editing by leveraging 2D diffusion model [40], such as InstructPix2Pix [4]. Subsequent methods [51, 9, 50, 34], inspired by IN2N, iteratively edit rendered images while updating the underlying 3D representation. However, these approaches independently edit single views with 2D diffusion models, neglecting multi-view consistency, as illustrated in Fig. 2-(a). More recent methods [11, 8, 7] incorporate multi-view information, but they lack a comprehensive strategy for consistent view alignment. For example, [11, 8] use only a limited set of key views, requiring additional diffusion passes or extra layers to integrate multi-view information. Similarly, [57] trains a diffusion model to achieve multi-view consistency, incurring high costs. Consequently, existing methods still exhibit multi-view inconsistency, resulting in noisy gradients that hinder optimization and produce suboptimal outputs, such as minimal edits or blur outputs.
@@
To tackle the limitations in both the multi-view inconsistency and the optimization inefficiency of pre-trained Gaussians in editing, we propose a novel 3D scene editing framework, EditSplat. To ensure multi-view consistency between edited images, we propose a Multi-view Fusion Guidance (MFG) method that ensures multi-view consistent editing utilizing a 2D diffusion model and the geometric properties of 3DGS. Inspired by [11, 7], our method projects initially edited multi-view images onto a target view using 3DGS depth maps, enabling smooth blending based on depth values across views to integrate multi-view information. Leveraging classifier-free guidance in the text-image diffusion process [16, 4], we carefully edit source images based on text prompt by incorporating multi-view fused images into the diffusion process, ensuring robust alignment of edited images with multi-view details. To further preserve source fidelity, we utilize the source images as auxiliary guidance. By balancing guidance scores between multi-view fused images, source images, and text prompts, our method enables the diffusion model to edit source images to the multi-view-consistent outputs, aligning with essential multi-view information. Unlike iterative editing approaches that require expensive optimization processes during rendered image editing, our proposed EditSplat edits source images directly, allowing for immediate 3DGS model optimization.
Figure 2: Challenging cases. (a) Standard 2D editing yields inconsistent bear textures, whereas Multi-view Fusion Guidance (MFG) achieves consistent 2D edits. (b) Rendered clown images without Attention-Guided Trimming (AGT) show less optimized areas (e.g., nose), while AGT enhances optimization in deeper red.
@@ 2 Related Work
Editing 3D scenes with texts aims to precisely and efficiently modify 3D representations. Recent advancements [35, 58, 28, 26] have used Score Distillation Sampling (SDS) for this task. Iterative methods such as InstructNerf2Nerf [12] and its extensions [9, 51, 50, 34] repeatedly edit rendered images while optimizing the underlying 3D representation. However, those works often overlook multi-view information, leading to multi-view inconsistency—a significant challenge. Certain approaches [8, 19] utilize the epipolar lines, though they still struggle to preserve fine-grained details.
Closely related to our work, VICA-NeRF [11] projects a limited set of key views to non-key views, resulting in noisy images and requiring two additional diffusion steps. ConsistDreamer [7] also projects all views inefficiently, leading to noisy images and necessitating extensive U-Net training.
In comparison, EditSplat efficiently integrates multi-view information through a carefully designed multi-view fusion strategy involving sequential projection and blending, avoiding extra training, diffusion processes, or additional layers [8] for multi-view consistent and fine-grained editing. 3DEgo  [23] employs classifier-free guidance with IP2P, similar to our approach. However, each edit for a single view requires guidance from all frames, leading to inconsistent edits due to misaligned viewpoints across frames. In contrast, EditSplat efficiently integrates and blends multi-view images aligned with viewpoints, enabling more accurate multi-view consistent editing.
Furthermore, most existing methods [9, 8, 51, 11, 23] rely on binary masks from the source image to localize the editing area, which constrains edits and overlooks semantic information when the target object is not clearly defined. Inspired by DreamEditor [58] that projects attention maps to Mesh, EditSplat addresses this localization limitation by explicitly assigning attention maps to 3D Gaussians.
To the best of our knowledge, no existing method addresses the optimization inefficiency of pre-trained 3DGS for editing. We demonstrate that pruning 3D Gaussians leads to more effective densification and optimization.
@@ 3 Method
We propose a novel 3D editing framework, EditSplat, carefully designed to achieve three key objectives: (i) multi-view consistent editing, (ii) efficient optimization, and (iii) semantically rich local editing. To meet these goals, we introduce Multi-view Fusion Guidance (MFG) and Attention-Guided Trimming (AGT). MFG incorporates essential multi-view information into the diffusion model to ensure consistency. In AGT, attention weights from the diffusion model are explicitly assigned to each Gaussian in 3D Gaussian Splatting (3DGS). This enables efficient convergence by pruning Gaussians with high attention weight and selectively optimizing them for semantic local editing. The overall pipeline is shown in Fig. 3. In the following sections, we provide a review of 3DGS and classifier-free guidance in Sec. 3.1, followed by our primary methods: MFG in Sec. 3.2 and AGT in Sec. 3.3.
@@ 3.2 Multi-View Fusion Guidance (MFG)
For the multi-view projection process, we select the top 5 adjacent views, prioritizing them based on their proximity to the target view, determined by camera position and orientation in world coordinates. To manage overlapping pixels across multiple views, we implement an iterative alpha blending strategy guided by depth values. Starting from the farthest pixel (i.e., highest depth value), pixel pairs are blended iteratively in decreasing depth order. The final pixel color is determined by aligning and unprojecting the multi-view images into 3D space using depth maps and camera parameters, followed by reprojecting onto the target view to resolve overlapping pixel contributions. This approach ensures that each target view captures comprehensive details shared across multiple views. To handle sparse background regions and achieve seamless integration, we refine these areas using SAM [25] to fill them with source content. This sequential multi-view fusion strategy enriches each target view with consistent and detailed information, reinforcing alignment and coherence across all views. The resulting multi-view fusion image $h_{M}$ for each target view serves as an essential input for subsequent editing processes, integrating comprehensive visual and geometric details shared across the views. Note that further details are provided in the supplementary material (Sec. B.1).
@@
To effectively integrate multi-view fusion information into the diffusion process and edit source image without additional training [7], extra layers [8], or repeated diffusion passes [11], we leverage classifier-free guidance. This allows the direct incorporation of multi-view fusion details into the diffusion process. According to Liu et al. [31], a conditional diffusion model can combine score estimates from multiple conditioning values. Building on this, we adapt classifier-free guidance, facilitating the implicit classifier $p_{\theta}$ to assign high probabilities to multi-view fused visual features, color consistency, and structural properties. This enables the edited images to maintain a strong alignment with multi-view fused images throughout the editing process.
@@
Our method efficiently directs the diffusion process toward generating outputs that achieve seamless multi-view consistency, including multi-view fusion images’ texture, color, geometric, and source images. By balancing guidance scores between $s_{M}$, source images $s_{S}$, and text prompts $s_{T}$, our approach enables the diffusion model to align edits with essential multi-view information and generate consistent outputs. Unlike iterative editing approaches [9, 12, 50, 51] that costly edit rendered images and require optimization of the 3D model, our approach directly edits the source image, ensuring precise alignment with multi-view information and the text prompt.
@@ 3.3 Attention-Guided Trimming (AGT)
where $S_{v,j}$ denotes a set of indices of all attention weights affected by the $j$-th Gaussian from the attention map for the $v$-th view. $\text{Softmax}\left(\frac{Q_{v}K^{T}}{\sqrt{q}}\right)_{s}\in[0,1]$ denotes an individual attention weight at $s$ from an attention map for the $v$-th view. By aggregating attention weight across $V$, this weighting mechanism ensures that each Gaussian reflects the overall importance of multi-view attention maps, emphasizing regions that require pronounced geometrical or visual changes during editing. Details on preparing the attention map are in the supplementary material (Sec. C.1).
@@ 4.1 Implementation Details
All experiments are conducted with fixed hyperparameters across all scenes and prompts, except for the number of epochs (ranging from 5 to 10). For a fair comparison with the baselines, the guidance scale of MFG is set to $s_{T}=7.5$, $s_{M}=1.0$, and $s_{S}=0.5$, aligning with those of the baselines using IP2P, such as GaussianEditor and DGE, which have guidance scales of $s_{T}=7.5$ and $s_{I}=1.5$. We matched the text guidance scale $s_{T}$ with the baselines and adjusted the image guidance scale such that the sum of $s_{M}$ and $s_{S}$ equals $s_{I}$ to ensure comparable conditions. In AGT, the threshold value for localization $w_{thres}$ is set to 0.1, and the pruning proportion $k$ is 0.15.
@@ 4.4 Quantitative Comparisons
Despite the subjective nature of 3D scene generation and editing, we adhere to established practices [50, 51, 9, 35, 13, 26, 8] by employing CLIP [38] for quantitative evaluation. Our comparison metrics include two types of CLIP scores. CLIP text-image directional similarity captures the alignment between changes in text captions and the corresponding changes between rendered and source images in CLIP space. Meanwhile, CLIP text-image similarity measures the alignment between the CLIP embeddings of the target text and the rendered images. In addition, we perform user study based on human preference. The user study was conducted with 100 participants via Amazon Mechanical Turk [1]. All of our metrics are evaluated under 8 scenes with 2 prompts each to ensure fair comparison and diverse feedback. As shown in Table 1, EditSplat achieves the highest scores across all categories, demonstrating superior performance. This demonstrates that EditSplat yields results that most closely align with the text prompt while achieving the highest human perceptual quality. Details are provided in the supplementary materials (Sec. D.2).
@@ 5 Conclusion
We introduce EditSplat, a robust 3D scene editing framework comprising Multi-view Fusion Guidance (MFG) and Attention-Guided Trimming (AGT). MFG addresses the critical challenge of achieving multi-view consistency, which often results in suboptimal outputs such as minimal edits or blurriness. MFG integrates essential multi-view information into the diffusion process, ensuring consistent edits across all views, consequently enabling editing of source images and direct optimization of the 3D Gaussian Splatting model, unlike iterative editing approaches. AGT enhances optimization efficiency and enables precise and semantic local editing by pruning pre-trained Gaussians and selectively optimizing those relevant for precise edits. Extensive evaluations of various complex real-world scenes demonstrate that EditSplat outperforms existing methods, setting a new benchmark for text-driven 3D scene editing.
@@ A.1 Extensive Results
Figure 8: Multi-View Fusion Guidance (MFG) Editing Process. MFG resolves multi-view inconsistency in the initially edited images through a multi-view fusion process that integrates information from multiple views to produce multi-view fused images. These fused images are subsequently incorporated into the diffusion process, ensuring consistent editing across all views.

### v2

@@ Abstract
Recent advancements in 3D editing have highlighted the potential of text-driven methods in real-time, user-friendly AR/VR applications. However, current methods rely on 2D diffusion models without adequately considering multi-view information, resulting in multi-view inconsistency. While 3D Gaussian Splatting (3DGS) significantly improves rendering quality and speed, its 3D editing process encounters difficulties with inefficient optimization, as pre-trained Gaussians retain excessive source information, hindering optimization. To address these limitations, we propose EditSplat, a novel text-driven 3D scene editing framework that integrates Multi-view Fusion Guidance (MFG) and Attention-Guided Trimming (AGT). Our MFG ensures multi-view consistency by incorporating essential multi-view information into the diffusion process, leveraging classifier-free guidance from the text-to-image diffusion model and the geometric structure inherent to 3DGS. Additionally, our AGT utilizes the explicit representation of 3DGS to selectively prune and optimize 3D Gaussians, enhancing optimization efficiency and enabling precise, semantically rich local editing. Through extensive qualitative and quantitative evaluations, EditSplat achieves state-of-the-art performance, establishing a new benchmark for text-driven 3D scene editing. Project website: https://kuai-lab.github.io/editsplat2024/
@@
Figure 1: Result of EditSplat. EditSplat enables flexible and high-quality editing of pre-trained 3D Gaussian Splatting models guided solely by textual instructions. Through its design focused on multi-view consistency, efficient optimization, and precise semantic local editing, Our approach demonstrates robust performance, producing realistic and fine-grained 3D scene modifications.
@@ 1 Introduction
InstructNeRF2NeRF (IN2N) [12] introduced the pipeline for text-driven 3D editing by leveraging 2D diffusion model [41], such as InstructPix2Pix [4]. Subsequent methods [52, 9, 51, 35], inspired by IN2N, iteratively edit rendered images while updating the underlying 3D representation. However, these approaches independently edit single views with 2D diffusion models, neglecting multi-view consistency, as illustrated in Fig. 2-(a). More recent methods [11, 8, 7] incorporate multi-view information, but they lack a comprehensive strategy for consistent view alignment. For example, [11, 8] use only a limited set of key views, requiring additional diffusion passes or extra-parameterized layers to integrate multi-view information. Similarly, [7] trains a diffusion model to achieve multi-view consistency, incurring high costs. Consequently, existing methods still exhibit multi-view inconsistency, resulting in noisy gradients that hinder optimization and produce suboptimal outputs, such as minimal edits or blur outputs.
@@
To tackle the limitations in both the multi-view inconsistency and the optimization inefficiency of pre-trained Gaussians in editing, we propose a novel 3D scene editing framework, EditSplat. To ensure multi-view consistency between edited images, we propose a Multi-view Fusion Guidance (MFG) method that ensures multi-view consistent editing utilizing a 2D diffusion model and the geometric structure inherent to 3DGS. Inspired by [11, 7], our method projects initially edited multi-view images onto a target view using 3DGS depth maps, enabling smooth blending based on depth values across views to integrate multi-view information. Leveraging classifier-free guidance in the text-image diffusion process [16, 4], we carefully edit source images based on text prompt by incorporating multi-view fused images into the diffusion process, ensuring robust alignment of edited images with multi-view details. To further preserve source fidelity, we utilize the source images as auxiliary guidance. By balancing guidance scores between multi-view fused images, source images, and text prompts, our method enables the diffusion model to edit source images to the multi-view-consistent outputs, aligning with essential multi-view information. Unlike iterative editing approaches that require expensive optimization processes during rendered image editing, our proposed EditSplat edits source images directly, allowing for immediate 3DGS model optimization.
Figure 2: Challenging cases. (a) Conventional 2D editing without Multi-view Fusion Guidance (MFG) results in inconsistent textures across different views (e.g., bear fur), whereas applying MFG achieves consistent multi-view edits. (b) Editing without Attention-Guided Trimming (AGT) leads to inefficient optimization, resulting in less edited regions (e.g., clown’s nose). AGT effectively enhances optimization quality, producing richer colors.
@@ 2 Related Work
Text-driven 3D scene editing aims to modify 3D representations precisely and efficiently using textual instructions. Despite recent progress, key challenges such as multi-view consistency, semantic localization, and optimization efficiency remain.
Recent advancements [36, 59, 28, 26] employ Score Distillation Sampling (SDS) [38], while iterative methods such as InstructNerf2Nerf [12] and its extensions [9, 52, 51, 35, 33] repeatedly edit rendered images while optimizing the underlying 3D representation. However, those works often overlook multi-view information, leading to inconsistency.
Recent methods explicitly addressing multi-view consistency, VICA-NeRF [11] and ConsistDreamer [7], employ view projection strategies but suffer from noisy projection outputs and computational overhead due to additional diffusion steps [11] or extensive U-Net [42] training [7]. Other approaches [8, 19] utilize the epipolar lines, though they still struggle to preserve fine-grained details. Moreover, 3DEgo [23] employs classifier-free guidance with IP2P [4], but its requirement for guidance from all frames in single-view edits inevitably results in inconsistencies due to misalignment viewpoint.
Furthermore, most existing methods [9, 8, 52, 11, 23] rely on binary masks from the source image to localize the editing area, which constrains edits and overlooks semantic information when the target object is not clearly defined.
For multi-view consistency, we introduce a Multi-view Fusion Guidance (MFG) involving sequential projection and blending, avoiding noise issues and the additional training [7] or diffusion refinement [11]. Regarding semantic localization, inspired by DreamEditor [59], EditSplat explicitly assigns attention maps to 3D Gaussians, achieving higher semantic precision without binary mask constraints. Additionally, as far as we know, we are the first to address the optimization inefficiency inherent in pre-trained 3DGS models by strategically pruning 3D Gaussians, enhancing both densification and optimization efficiency.
@@ 3 Method
We propose a novel 3D editing framework, EditSplat, carefully designed to achieve three key objectives: (i) multi-view consistent editing, (ii) efficient optimization, and (iii) semantically rich local editing. To meet these goals, we introduce Multi-view Fusion Guidance (MFG) and Attention-Guided Trimming (AGT). MFG incorporates essential multi-view information into the diffusion model to ensure multi-view consistency. In AGT, attention weights from the diffusion model are explicitly assigned to each Gaussian in 3D Gaussian Splatting (3DGS). This enables efficient convergence by pruning Gaussians with high attention weight and selectively optimizing them for semantic local editing. The overall pipeline is shown in Fig. 3. In the following sections, we provide a review of 3DGS and classifier-free guidance in Sec. 3.1, followed by our primary methods: MFG in Sec. 3.2 and AGT in Sec. 3.3.
@@ 3.2 Multi-View Fusion Guidance (MFG)
For the multi-view projection process, we select the top 5 adjacent views, prioritizing them based on their proximity to the target view, determined by camera position and orientation in world coordinates. To manage overlapping pixels across multiple views, we implement an iterative alpha blending strategy guided by depth values. Starting from the farthest pixel (i.e., highest depth value), pixel pairs are blended iteratively in decreasing depth order. The final pixel color is determined by aligning and unprojecting the multi-view images into 3D space using depth maps and camera parameters, followed by reprojecting onto the target view to resolve overlapping pixel contributions. This approach ensures that each target view captures comprehensive details shared across multiple views. To handle sparse background regions and achieve seamless integration, we refine these areas using SAM [25] to fill them with source content. This sequential multi-view fusion strategy enriches each target view with consistent and detailed information, reinforcing alignment and coherence across all views. The resulting multi-view fusion image $h_{M}$ for each target view serves as an essential input for subsequent editing processes, integrating comprehensive visual and geometric details shared across the views. Note that further details are provided in the supplementary material.
@@
To effectively integrate multi-view fusion information into the diffusion process and edit source images without additional training [7], extra-parameterized layers [8], or repeated diffusion passes [11], we leverage classifier-free guidance. This allows the direct incorporation of multi-view fusion details into the diffusion process. According to Liu et al. [31], a conditional diffusion model can combine score estimates from multiple conditioning values. Building on this, we adapt this approach, facilitating the implicit classifier $p_{\theta}$ to assign high probabilities to multi-view fused visual features, color consistency, and structural properties. This ensures the edited images maintain a strong alignment with multi-view fused images throughout the editing process.
@@
Our method efficiently directs the diffusion process toward generating outputs that achieve seamless multi-view consistency, including multi-view fusion images’ texture, color, geometric, and source images. By balancing guidance scores between $s_{M}$, source images $s_{S}$, and text prompts $s_{T}$, our approach enables the diffusion model to align edits with essential multi-view information and generate consistent outputs. Unlike iterative editing approaches [9, 12, 51, 52] that iteratively edit rendered images while updating the underlying 3D representation, our approach directly edits the source image, ensuring precise alignment with multi-view information and the text prompt.
@@ 3.3 Attention-Guided Trimming (AGT)
where $S_{v,j}$ denotes a set of indices of all attention weights affected by the $j$-th Gaussian from the attention map for the $v$-th view. $\text{Softmax}\left(\frac{Q_{v}K^{T}}{\sqrt{q}}\right)_{s}\in[0,1]$ denotes an individual attention weight at $s$ from an attention map for the $v$-th view. By aggregating attention weight across $V$, this weighting mechanism ensures that each Gaussian reflects the overall importance of multi-view attention maps, emphasizing regions that require pronounced geometrical or visual changes during editing. Details on preparing the attention map are in the supplementary material.
@@ 4.1 Implementation Details
All experiments are conducted with fixed hyperparameters across all scenes and prompts, except for the number of epochs (ranging from 5 to 10). For a fair comparison with baseline methods such as GaussianEditor [9] and DGE [8], which adopt the IP2P [4] and set to $s_{T}=7.5$, $s_{I}=1.5$, we set the guidance scales of MFG to $s_{T}=7.5$, $s_{M}=1.0$, and $s_{S}=0.5$. Specifically, we matched the text guidance scale $s_{T}$ with the baselines and adjusted the image guidance scales such that the sum of $s_{M}$ and $s_{S}$ equals $s_{I}$, ensuring comparable experimental conditions. In AGT, the threshold value for localization $w_{thres}$ is set to $0.1$, and the pruning proportion $k$ is $0.15$.
@@ 4.4 Quantitative Comparisons
Despite the subjective nature of 3D scene generation and editing, we adhere to established practices [51, 52, 9, 36, 13, 26, 8] by employing CLIP [39] for quantitative evaluation. Our comparison metrics include two types of CLIP scores. CLIP text-image directional similarity captures the alignment between changes in text captions and the corresponding changes between rendered and source images in CLIP space. Meanwhile, CLIP text-image similarity measures the alignment between the CLIP embeddings of the target text and the rendered images. In addition, we perform user study based on human preference. The user study was conducted with 100 participants via Amazon Mechanical Turk [1]. All of our metrics are evaluated under 8 scenes with 2 prompts each to ensure fair comparison and diverse feedback. As shown in Table 1, EditSplat achieves the highest scores across all categories, demonstrating superior performance. This demonstrates that EditSplat yields results that most closely align with the text prompt while achieving the highest human perceptual quality. Details are provided in the supplementary materials.
@@ 5 Conclusion
We introduce EditSplat, a robust 3D scene editing framework comprising Multi-view Fusion Guidance (MFG) and Attention-Guided Trimming (AGT). MFG addresses the critical challenge of achieving multi-view consistency, which often results in suboptimal outputs such as minimal edits or blurred results. By integrating critical multi-view information into the diffusion process, MFG ensures consistent and precise edits across different viewpoints, facilitating editing of source images and direct optimization of the 3D Gaussian Splatting model, unlike iterative editing approaches. AGT enhances optimization efficiency and enables precise and semantically rich local editing by pruning pre-trained Gaussians and selectively optimizing those relevant. Extensive evaluations of various complex real-world scenes demonstrate that EditSplat achieves state-of-the-art performance existing methods, setting a new benchmark for text-driven 3D scene editing.
@@
Acknowledgements
This research was supported by Korea Research Institute for defense Technology planning and advancement - Grant funded by Defense Acquisition Program Administration(DAPA)((KRIT-CT-23-021), 90%), the Ministry of Science and ICT (MSIT) of Korea, under the National Research Foundation (NRF) grant (RS-2024-00337548, 5%; RS-2025-00521602, 1%), the Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2024 (International Collaborative Research and Global Talent Development for the Development of Copyright Management and Protection Technologies for Generative AI, RS-2024-00345025, 3%), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2019-II190079, Artificial Intelligence Graduate School Program(Korea University), 1%), and Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT(MSIT, Korea)&Gwangju Metropolitan City.
@@ A.1 Extensive Results
Figure 8: Multi-View Fusion Guidance (MFG) 2D Editing Process. MFG resolves multi-view inconsistency in the initially edited images through a multi-view fusion process that integrates information from multiple views to produce multi-view fused images. These fused images are subsequently incorporated into the diffusion process, ensuring consistent editing across all views.